# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ Abstract
Concept-based Models are neural networks that learn a concept extractor to map inputs to high-level concepts and an inference layer to translate these into predictions. Ensuring these modules produce interpretable concepts and behave reliably in out-of-distribution is crucial, yet the conditions for achieving this remain unclear. We study this problem by establishing a novel connection between Concept-based Models and reasoning shortcuts (RSs), a common issue where models achieve high accuracy by learning low-quality concepts, even when the inference layer is fixed and provided upfront. Specifically, we first extend RSs to the more complex setting of Concept-based Models and then derive theoretical conditions for identifying both the concepts and the inference layer. Our empirical results highlight the impact of reasoning shortcuts and show that existing methods, even when combined with multiple natural mitigation strategies, often fail to meet these conditions in practice.
@@ 1 Introduction
Concept-based Models (CBMs) are a broad class of self-explainable classifiers [Alvarez-Melis and Jaakkola, 2018, Koh et al., 2020, Espinosa Zarlenga et al., 2022, Marconato et al., 2022, Taeb et al., 2022] designed for high performance and ante-hoc interpretability. Learning a CBM involves solving two conjoint problems: acquiring high-level concepts describing the input (e.g., an image) and an inference layer that predicts a label from them. In many applications, it is essential that these two elements are “high quality”, in the sense that: i) the concepts should be interpretable, as failure in doing so compromises understanding [Schwalbe, 2022, Poeta et al., 2023] and steerability [Teso et al., 2023, Gupta and Narayanan, 2024], both key selling point of CBMs; and ii) the concepts and inference layer should behave well also out of distribution (OOD), e.g., they should not pick up spurious correlations between the input, the concepts and the output [Geirhos et al., 2020, Bahadori and Heckerman, 2021, Stammer et al., 2021].
Figure 1: Joint reasoning shortcuts. The goal is to predict whether the sum of two MNIST digits is odd (as in Example 1) from a training set of all possible unique (even, even), (odd, odd), and (odd, even) pairs of MNIST digits. Green elements are fixed, purple ones are learned. Left: ground-truth concepts and inference layer. Middle: NeSy-CBMs with given knowledge can learn reasoning shortcuts, i.e., concepts with unintended semantics. Right: CBMs can learn joint reasoning shortcuts, i.e., both concepts and inference layer have unintended semantics.
Mighty Jesus please love me as much as I love you.
@@
It is natural to ask under what conditions CBMs can acquire such “high-quality” concepts and inference layers. While existing studies focus on concept quality [Mahinpei et al., 2021, Zarlenga et al., 2023, Marconato et al., 2023a], they neglect the role of the inference layer altogether. In contrast, we cast the question in terms of whether it is possible to identify from data concepts and inference layers with the intended semantics, defined formally in Section 3.1. We proceed to answer this question by building a novel connection with reasoning shortcuts (RSs), a well-known issue in Neuro-Symbolic (NeSy) AI whereby models achieve high accuracy by learning low-quality concepts even if the inference layer is fixed [Chang et al., 2020, Marconato et al., 2023b, Wang et al., 2023, Yang et al., 2024a]. E.g., a NeSy model for autonomous driving whose inference layer encodes the traffic laws might confuse pedestrians with red lights, as both entail the same prediction (the car has to stop) [Bortolotti et al., 2024].
We generalize RSs to CBMs in which the inference layer is learned, and supervision on the concepts may be absent. Our analysis shows that shortcut solutions can be exponentially many (indeed, we count them explicitly in Section B.2) and that, in general, maximum likelihood training is insufficient for attaining intended semantics. This negatively affects interpretability – as the learned concepts cannot be easily mapped to the ground-truth annotations – and OOD behavior. On the positive side, we also specify conditions under which (under suitable assumptions) CBMs cannot be fooled by RSs, proving that the ground-truth concepts and inference layer can be identified (see Theorem 3.8).
@@
Our numerical simulations with regular and NeSy CBMs on several learning tasks suggest CBMs can be severely impacted by reasoning shortcuts in practice, as expected, and also that unfortunately the benefits of popular mitigation strategies do not carry over to this more challenging problem. These results cast doubts on the ability of these models to identify concepts and inference layers with the intended semantics unless appropriately nudged.
Contributions: In summary, we:
  • Generalize reasoning shortcuts to the challenging case of CBMs whose inference layer is learned.
  • Study conditions for maximum likelihood training to identify high quality concepts and inference layers.
  • Present empirical evidence that proven mitigations fail in this challenging problem.
@@ 2 Preliminaries
Concept-based Models (CBMs) first map the input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ into $k$ discrete concepts $\bm{\mathrm{c}}=(c_{1},\ldots,c_{k})\in\mathcal{C}$ via a neural backbone, and then infer labels $\bm{\mathrm{y}}\in\mathcal{Y}$ from this using a white-box layer, e.g., a linear layer. This setup makes it easy to figure out what concepts are most responsible for any prediction, yielding a form of ante-hoc concept-based interpretability. Several architectures follow this recipe, including approaches for converting black-box neural network classifiers into CBMs [Yuksekgonul et al., 2022, Wang et al., 2024, Dominici et al., 2024, Marcinkevičs et al., 2024].
A key issue is how to ensure the concepts are interpretable. Some CBMs rely on concept annotations [Koh et al., 2020, Sawada and Nakamura, 2022, Espinosa Zarlenga et al., 2022, Marconato et al., 2022, Kim et al., 2023, Debot et al., 2024]. These are however expensive to obtain, prompting researchers to replace them with (potentially unreliable [Huang et al., 2024, Sun et al., 2024]) annotations obtained from foundation models [Oikarinen et al., 2022, Yang et al., 2023, Srivastava et al., 2024] or unsupervised concept discovery [Alvarez-Melis and Jaakkola, 2018, Chen et al., 2019, Taeb et al., 2022, Schrodi et al., 2024].
Neuro-Symbolic CBMs (NeSy-CBMs) specialize CBMs to tasks in which the prediction $\bm{\mathrm{y}}\in\mathcal{Y}$ ought to comply with known safety or structural constraints. These are supplied as a formal specification – a logic formula $\mathsf{K}$, aka knowledge – tying together the prediction $\bm{\mathrm{y}}$ and the concepts $\bm{\mathrm{c}}$. In NeSy-CBMs, the inference step is a symbolic reasoning layer that steers [Lippi and Frasconi, 2009, Diligenti et al., 2012, Donadello et al., 2017, Xu et al., 2018] or guarantees [Manhaeve et al., 2018, Giunchiglia and Lukasiewicz, 2020, Hoernle et al., 2022, Ahmed et al., 2022] the labels and concepts to be logically consistent according to $\mathsf{K}$. Throughout, we will consider the following example task:
Example 1 (MNIST-SumParity).
Given two MNIST digits [LeCun, 1998], we wish to predict whether their sum is even or odd. The numerical values of the two digits can be modelled as concepts $\bm{\mathrm{C}}\in\{0,\ldots,9\}^{2}$, and the inference layer is entirely determined by the prior knowledge: $\mathsf{K}=((y=1)\Leftrightarrow\text{$(C_{1}+C_{2})$ is odd})$. This specifies that the label $y\in\{0,1\}$ ought to be consistent with the predicted concepts $\bm{\mathrm{C}}$. See Fig. 1 for an illustration.
Like CBMs, NeSy-CBMs are usually trained via maximum likelihood and gradient descent, but without concept supervision. The reasoning layer is typically imbued with fuzzy [Zadeh, 1988] or probabilistic [De Raedt and Kimmig, 2015] logic semantics to ensure differentiability. Many NeSy-CBMs require $\mathsf{K}$ to be provided upfront, as in Example 1, hence their inference layer has no learnable parameters. Starting with Section 3, we will instead consider NeSy-CBMs that – just like regular CBMs – learn the inference layer [Wang et al., 2019, Liu et al., 2023, Daniele et al., 2023, Tang and Ellis, 2023, Wüst et al., 2024].
@@ 2.1 Reasoning Shortcuts
Before discussing reasoning shortcuts, we need to establish a clear relationship between concepts, inputs, and labels. The RS literature does so by assuming the following data generation process [Marconato et al., 2023b, Yang et al., 2024a, Umili et al., 2024]: each input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ is the result of sampling $k$ ground-truth concepts $\bm{\mathrm{g}}=(g_{1},\ldots,g_{k})\in\mathcal{G}$ (e.g., in MNIST-SumParity two numerical digits) from an unobserved distribution $p^{*}(\bm{\mathrm{G}})$ and then $\bm{\mathrm{x}}$ itself (e.g., two corresponding MNIST images) from the conditional distribution $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$. [^1] This distribution subsumes stylistic factors, e.g., calligraphy. Labels $\bm{\mathrm{y}}\in\mathcal{Y}$ are sampled from the conditional distribution of $\bm{\mathrm{g}}$ given by $p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{g}};\mathsf{K})$ consistently with $\mathsf{K}$ (e.g., $y=1$ if and only if $g_{1}+g_{2}$ is odd). The ground-truth distribution is thus:
@@
Intuitively, a reasoning shortcut (RS) occurs when a NeSy-CBM with fixed knowledge $\mathsf{K}$ attains high or even perfect label accuracy by learning concepts $\bm{\mathrm{C}}$ that differ from the ground-truth ones $\bm{\mathrm{G}}$ [Marconato et al., 2023b]. As commonly done, we work in the setting where $\mathcal{G}=\mathcal{C}$.
Example 2.
[…]
RSs by definition yield good labels in-distribution, yet they compromise out-of-distribution (OOD) performance. For instance, in autonomous driving tasks, NeSy-CBMs can confuse the concepts of “pedestrian” and “red light”, leading to poor decisions for OOD decisions where the distinction matters [Bortolotti et al., 2024]. The meaning of concepts affected by RSs is unclear, affecting understanding [Schwalbe, 2022], intervenability [Shin et al., 2023, Espinosa Zarlenga et al., 2024, Steinmann et al., 2024b], debugging [Lertvittayakumjorn et al., 2020, Stammer et al., 2021] and down-stream applications that hinge on concepts being high-quality, like NeSy formal verification [Xie et al., 2022, Zaid et al., 2023, Morettin et al., 2024]. Unfortunately, existing works on RSs do not apply to CBMs and NeSy-CBMs where the inference layer is learned.
Figure 2: Examples of semantics in MNIST-SumParity restricted to $\bm{\mathrm{g}},\bm{\mathrm{c}}\in\{0,1,2\}^{2}$ for readability. Left: ideally, $\bm{\mathrm{\alpha}}$ should be the identity (i.e., $\bm{\mathrm{C}}$ recovers the ground-truth concepts $\bm{\mathrm{G}}$) and the inference layer should learn $\bm{\mathrm{\beta}}^{*}$. Middle: $(\bm{\mathrm{\alpha}}_{\mathrm{IS}},\bm{\mathrm{\beta}}_{\mathrm{IS}})\neq(\mathrm{id},\bm{\mathrm{\beta}}^{*})$ has intended semantics (Definition 3.3), i.e., the ground-truth concepts and inference layer can be easily recovered and generalize OOD. Right: $(\bm{\mathrm{\alpha}}_{\mathrm{JRS}},\bm{\mathrm{\beta}}_{\mathrm{JRS}})$ affected by the JRS in Fig. 1. Elements (predicted concepts $\bm{\mathrm{C}}$ and entries in $\bm{\mathrm{\beta}}$) in red are never predicted nor used, highlighting simplicity bias. Maps are visualized as matrices.
@@ 3 Reasoning Shortcuts in CBMs
Given a finite set $\mathcal{S}$, we indicate with $\Delta_{\mathcal{S}}\subset[0,1]^{|\mathcal{S}|}$ the simplex of probability distributions $P(S)$ over items in $\mathcal{S}$. Notice that the set of its vertices $\mathsf{Vert}(\Delta_{\mathcal{S}})$ contains all point mass distributions $\mathbbm{1}\!\{S=s\}$ for all $s\in\mathcal{S}$.
CBMs as pairs of functions. CBMs and NeSy-CBMs differ in how they implement the inference layer, hence to bridge them we employ the following unified formalism. Any CBM can be viewed as a pair of learnable functions implementing the concept extractor and the inference layer, respectively, cf. Fig. 2. Formally, the former is a function $\bm{\mathrm{f}}:\mathbb{R}^{n}\to\Delta_{\mathcal{C}}$ mapping inputs $\bm{\mathrm{x}}$ to a conditional distribution $p(\bm{\mathrm{C}}\mid\bm{\mathrm{x}})$ over the concepts, however it can be better understood as a function $\bm{\mathrm{\alpha}}:\mathcal{G}\to\Delta_{\mathcal{C}}$ taking ground-truth concepts $\bm{\mathrm{g}}$ as input instead, and defined as:
On the other hand, the inference layer is a function $\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$ mapping the concept distribution output by the concept extractor into a label distribution $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{f}}(\bm{\mathrm{x}}))$. For clarity, we also define $\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$, which is identical to the former except it works with concept values rather than distributions, that is:
@@
Hence, a CBM entails both a pair $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ and a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ [^2] We work in the non-parametric setting, hence $\mathcal{F}$ and $\Omega$ contain all learnable concept extractors $\bm{\mathrm{f}}$ and inference layers $\bm{\mathrm{\omega}}$, and similarly $\mathcal{A}$ and $\mathcal{B}$ contain all learnable maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. Later on, we will make use of the fact that $\mathcal{A}$ and $\mathcal{B}$ are simplices [Morton, 2013, Montúfar et al., 2014], i.e., each $\bm{\mathrm{\alpha}}\in\mathcal{A}$ (resp. $\bm{\mathrm{\beta}}\in\mathcal{B}$) can be written as a convex combination of vertices $\mathsf{Vert}(\mathcal{A})$ (resp. $\mathsf{Vert}(\mathcal{B})$).
As mentioned in Section 2, supplying prior knowledge $\mathsf{K}$ to a NeSy-CBM is equivalent to fixing the inference layer to a corresponding function $\bm{\mathrm{\omega}}^{*}$ (and $\bm{\mathrm{\beta}}^{*}$). Note that whereas $\bm{\mathrm{\omega}}^{*}$ changes based on how reasoning is implemented – e.g., fuzzy vs. probabilistic logic – $\bm{\mathrm{\beta}}^{*}$ does not, as both kinds of reasoning layers behave identically when input any point-mass concept distribution $\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\}$.
Standard assumptions. The maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$ are especially useful for analysis provided the following two standard assumptions about how data are distributed:
Assumption 3.1 (Invertibility).
@@
This means that the ground-truth concepts $\bm{\mathrm{g}}$ can always be unambiguously recovered for all inputs $\bm{\mathrm{x}}$ by a sufficiently expressive concept extractor; the $\bm{\mathrm{\alpha}}$ it induces is the identity, indicated as $\mathrm{id}(\bm{\mathrm{g}}):=\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{g}}\}\in\mathsf{Vert}(\mathcal{A})$.
@@
This ensures that the labels $\bm{\mathrm{y}}$ can be predicted without any ambiguity from the ground-truth concepts $\bm{\mathrm{g}}$. Both assumptions apply in several tasks, see [Bortolotti et al., 2024]. The maximum log-likelihood objective is then written as:
@@
Here, $\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}$ is the conditional probability of the ground-truth labels $\bm{\mathrm{y}}$. Notice that under the above assumptions CBMs attaining maximum likelihood perfectly model the ground-truth data distribution $p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})$, see Lemma C.1.

### v2

@@ Abstract
Concept-based Models are neural networks that learn a concept extractor to map inputs to high-level concepts and an inference layer to translate these into predictions. Ensuring these modules produce interpretable concepts and behave reliably in out-of-distribution is crucial, yet the conditions for achieving this remain unclear. We study this problem by establishing a novel connection between Concept-based Models and reasoning shortcuts (RSs), a common issue where models achieve high accuracy by learning low-quality concepts, even when the inference layer is fixed and provided upfront. Specifically, we extend RSs to the more complex setting of Concept-based Models and derive theoretical conditions for identifying both the concepts and the inference layer. Our empirical results highlight the impact of RSs and show that existing methods, even combined with multiple natural mitigation strategies, often fail to meet these conditions in practice.
@@ 1 Introduction
Concept-based Models (CBMs) are a broad class of self-explainable classifiers [1, 2, 3, 4, 5, 6, 7] designed for high performance and ante-hoc interpretability. Learning a CBM involves solving two conjoint problems: acquiring high-level concepts describing the input (e.g., an image) and an inference layer that predicts a label from them. In many applications, it is essential that these two elements are “high quality”, in the sense that: i) the concepts should be interpretable, as failure in doing so compromises understanding [8, 9] and steerability [10, 11], both key selling points of CBMs; and ii) the concepts and inference layer should behave well also out of distribution (OOD), e.g., they should not pick up spurious correlations between the input, the concepts and the output [12, 13, 14].
@@
This raises the question of when CBMs can acquire “high-quality” concepts and inference layers. While existing studies focus on concept quality [15, 16, 17], they neglect the role of the inference layer altogether. In contrast, we cast the question in terms of whether it is possible to identify from data concepts and inference layers with the intended semantics, defined formally in Section˜3.1. We proceed to answer this question by building a novel connection with reasoning shortcuts (RSs), a well-known issue in Neuro-Symbolic (NeSy) AI whereby models achieve high accuracy by learning low-quality concepts even if the inference layer is fixed [18, 19, 20, 21]. For instance, a NeSy model for autonomous driving whose inference layer encodes the traffic laws might confuse pedestrians with red lights, as both entail the same prediction (the car has to stop) [22]. We generalize RSs to CBMs in which the inference layer is learned, and concept supervision may be absent. Our analysis shows that shortcuts can be exponentially many, even more than RSs, (we count them explicitly in Section˜B.5) and that maximum likelihood training is insufficient for attaining intended semantics. This hinders both interpretability and OOD behavior. On the positive side, we also specify conditions under which (under suitable assumptions) CBMs cannot be fooled by RSs, proving that the ground-truth concepts and inference layer can be identified (see Theorem˜3.9).
@@
Our evaluation on several learning tasks suggest that CBMs can be severely impacted by reasoning shortcuts in practice, as expected, and also that the benefits of popular mitigation strategies do not carry over to this more challenging problem. These results cast doubts on the ability of these models to identify concepts and inference layers with the intended semantics unless appropriately nudged.
Contributions: In summary, we: (i) generalize reasoning shortcuts to the challenging case of CBMs whose inference layer is learned (Section˜3.1), (ii) study conditions under which maximum likelihood training can identify good concepts and inference layers (Section˜3.2), and (iii) present empirical evidence that well-tested mitigations fail in this challenging setting (Section˜5).
Figure 1: Joint reasoning shortcuts. The goal is to predict whether the sum of two MNIST digits is odd (as in Example˜2.1) from a training set of all possible unique (even, even), (odd, odd), and (odd, even) pairs of MNIST digits. Green elements are fixed, purple ones are learned. Left: ground-truth concepts and inference layer. Middle: NeSy-CBMs with given knowledge can learn reasoning shortcuts, i.e., concepts with unintended semantics. Right: CBMs can learn joint reasoning shortcuts, i.e., both concepts and inference layer have unintended semantics.
@@ 2 Preliminaries
Concept-based Models (CBMs) first map the input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ into $k$ discrete categorical concepts $\bm{\mathrm{c}}=(c_{1},\ldots,c_{k})\in\mathcal{C}$ via a neural backbone, and then infer labels $\bm{\mathrm{y}}\in\mathcal{Y}$ from this using a white-box layer, e.g., a linear layer. This setup makes it easy to figure out what concepts are most responsible for any prediction, yielding a form of ante-hoc concept-based interpretability. Several architectures follow this recipe, including approaches for converting black-box neural networks into CBMs [23, 24, 25, 26].
A key issue is how to ensure the concepts are interpretable. Some CBMs rely on concept annotations [3, 27, 4, 5, 28, 29]. These are however expensive to obtain, prompting researchers to replace them with (potentially unreliable [30, 31]) annotations obtained from foundation models [32, 33, 34, 35] or unsupervised concept discovery [1, 2, 6, 36].
Neuro-Symbolic CBMs (NeSy-CBMs) specialize CBMs to tasks in which the prediction $\bm{\mathrm{y}}\in\mathcal{Y}$ ought to comply with known safety or structural constraints. These are supplied as a formal specification – a logic formula $\mathsf{K}$, aka knowledge – tying together the prediction $\bm{\mathrm{y}}$ and the concepts $\bm{\mathrm{c}}$. In NeSy-CBMs, the inference step is a symbolic reasoning layer that steers [37, 38, 39] or guarantees [40, 41, 42, 43, 44] the labels and concepts to be logically consistent according to $\mathsf{K}$. Throughout, we will consider this example task:
Example 2.1 (MNIST-SumParity).
Given two MNIST digits [45], we wish to predict whether their sum is even or odd. The numerical values of the two digits can be modelled as concepts $\bm{\mathrm{C}}\in\{0,\ldots,9\}^{2}$, and the inference layer is entirely determined by the prior knowledge: $\mathsf{K}=((y=1)\Leftrightarrow\text{$(C_{1}+C_{2})$ is odd})$. This specifies that the label $y\in\{0,1\}$ ought to be consistent with the predicted concepts $\bm{\mathrm{C}}$. See Fig.˜1 for an illustration.
Like CBMs, NeSy-CBMs are usually trained via maximum likelihood and gradient descent, but without concept supervision. The reasoning layer is typically imbued with fuzzy [46] or probabilistic [47] logic semantics to ensure differentiability. Many NeSy-CBMs require $\mathsf{K}$ to be provided upfront, as in Example˜2.1, hence their inference layer has no learnable parameters. Starting with Section˜3, we will instead consider NeSy-CBMs that – just like regular CBMs – learn the inference layer [48, 49, 50, 51, 52].
@@ 2.1 Reasoning Shortcuts
Before discussing reasoning shortcuts, we need to establish a clear relationship between concepts, inputs, and labels. The RS literature does so by assuming the following data generation process [19, 21, 53]: each input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ is the result of sampling $k$ ground-truth concepts $\bm{\mathrm{g}}=(g_{1},\ldots,g_{k})\in\mathcal{G}$ (e.g., in MNIST-SumParity two numerical digits) from an unobserved distribution $p^{*}(\bm{\mathrm{G}})$ and then $\bm{\mathrm{x}}$ itself (e.g., two corresponding MNIST images) from the conditional distribution $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$. [^1] This distribution subsumes stylistic factors, e.g., calligraphy. Labels $\bm{\mathrm{y}}\in\mathcal{Y}$ are sampled from the conditional distribution of $\bm{\mathrm{g}}$ given by $p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{g}};\mathsf{K})$ consistently with $\mathsf{K}$ (e.g., $y=1$ if and only if $g_{1}+g_{2}$ is odd). The ground-truth distribution is thus:
@@
Intuitively, a reasoning shortcut (RS) occurs when a NeSy-CBM with fixed knowledge $\mathsf{K}$ attains high or even perfect label accuracy by learning concepts $\bm{\mathrm{C}}$ that differ from the ground-truth ones $\bm{\mathrm{G}}$.
Example 2.2.
In MNIST-SumParity, a NeSy model can achieve perfect accuracy by mapping each pair of MNIST images $\bm{\mathrm{x}}=(\bm{\mathrm{x}}_{1},\bm{\mathrm{x}}_{2})$ to the corresponding ground-truth digits, that is, $\bm{\mathrm{c}}=(g_{1},g_{2})$. However, it would achieve the same accuracy if it were to map it to $\bm{\mathrm{c}}=(g_{1}\ \mathrm{mod}\ 2,g_{2}\ \mathrm{mod}\ 2)$, as doing so leaves the parity of the sum unchanged, see Fig.˜1. Hence, a NeSy model cannot distinguish between the two based on label likelihood alone during training.
RSs by definition yield good labels in-distribution, yet they compromise out-of-distribution (OOD) performance. For instance, in autonomous driving tasks, NeSy-CBMs can confuse the concepts of “pedestrian” and “red light”, leading to poor decisions for OOD decisions where the distinction matters [22]. The meaning of concepts affected by RSs is unclear, affecting understanding [8], intervenability [54, 55, 56], debugging [57, 14] and down-stream applications that hinge on concepts being high-quality, like NeSy formal verification [58, 59, 60]. Unfortunately, existing works on RSs do not apply to CBMs and NeSy-CBMs where the inference layer is learned.
As commonly done, we work in the setting with equal discrete predicted $\mathcal{C}$ and ground-truth $\mathcal{G}$ concept sets, i.e., $\mathcal{G}=\mathcal{C}$ [19, 21]. Notice that, we make no assumption on how the set $\mathcal{G}$ is made; multiple concept vocabularies at different levels of abstraction may be valid for a given task. At this stage, different choices of $\mathcal{G}$ are allowed but may lead to distinct results for RSs, depending on the number of ground-truth concepts and how they are related to the labels. Both these two aspects will be made clear in light of the data generation process as per Assumptions˜3.1 and 3.2.
Figure 2: Examples of semantics in MNIST-SumParity restricted to $\bm{\mathrm{g}},\bm{\mathrm{c}}\in\{0,1,2\}^{2}$ for readability. Left: ideally, $\bm{\mathrm{\alpha}}$ should be the identity (i.e., $\bm{\mathrm{C}}$ recovers the ground-truth concepts $\bm{\mathrm{G}}$) and the inference layer should learn $\bm{\mathrm{\beta}}^{*}$. Middle: $(\bm{\mathrm{\alpha}}_{\mathrm{IS}},\bm{\mathrm{\beta}}_{\mathrm{IS}})\neq(\mathrm{id},\bm{\mathrm{\beta}}^{*})$ has intended semantics (Definition˜3.3), i.e., the ground-truth concepts and inference layer can be recovered and generalize OOD. Right: $(\bm{\mathrm{\alpha}}_{\mathrm{JRS}},\bm{\mathrm{\beta}}_{\mathrm{JRS}})$ affected by the Joint Reasoning Shortcuts in Fig.˜1. Elements (predicted concepts $\bm{\mathrm{C}}$ and entries in $\bm{\mathrm{\beta}}$) in red are never predicted nor used, highlighting simplicity bias. Maps are visualized as matrices.
@@ 3 Reasoning Shortcuts in CBMs
Given a finite set $\mathcal{S}$, we indicate with $\Delta_{\mathcal{S}}\subset[0,1]^{|\mathcal{S}|}$ the simplex of probability distributions $P(Q)$ over items in $\mathcal{S}$. Any random variable $Q\in\mathcal{S}$ defines a point in the simplex via its distribution $P(Q)$. Notice that the set of the simplex vertices $\mathsf{Vert}(\Delta_{\mathcal{S}})$ contains all point mass distributions $\mathbbm{1}\!\{Q=q\}$ for all $q\in\mathcal{S}$. All relevant notation we will use is reported in Table˜6.
CBMs as pairs of functions. CBMs and NeSy-CBMs differ in how they implement the inference layer, hence to bridge them we employ the following unified formalism. Any CBM can be viewed as a pair of learnable functions implementing the concept extractor and the inference layer, respectively, cf. Fig.˜2. Formally, the former is a function $\bm{\mathrm{f}}:\mathbb{R}^{n}\to\Delta_{\mathcal{C}}$ mapping inputs $\bm{\mathrm{x}}$ to a conditional distribution $p(\bm{\mathrm{C}}\mid\bm{\mathrm{x}})$ over the concepts, however it can be better understood as a function $\bm{\mathrm{\alpha}}:\mathcal{G}\to\Delta_{\mathcal{C}}$ taking ground-truth concepts $\bm{\mathrm{g}}$ as input instead, and defined as:
$\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}):=\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{f}}(\bm{\mathrm{x}})]$ (2)
In contrast, the inference layer is a function $\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$ mapping the concept distribution output by the concept extractor into a label distribution $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{f}}(\bm{\mathrm{x}}))$. For clarity, we also define $\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$, which is identical to the former except it works with concept values rather than distributions, that is:
@@
Hence, a CBM entails both a pair $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ and a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$. [^2] We work in the non-parametric setting, hence $\mathcal{F}$ and $\Omega$ contain all learnable concept extractors $\bm{\mathrm{f}}$ and inference layers $\bm{\mathrm{\omega}}$, and similarly $\mathcal{A}$ and $\mathcal{B}$ contain all learnable maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. Later on, we will make use of the fact that $\mathcal{A}$ and $\mathcal{B}$ are simplices [61, 62], i.e., each $\bm{\mathrm{\alpha}}\in\mathcal{A}$ (resp. $\bm{\mathrm{\beta}}\in\mathcal{B}$) can be written as a convex combination of vertices $\mathsf{Vert}(\mathcal{A})$ (resp. $\mathsf{Vert}(\mathcal{B})$).
As mentioned in Section˜2, supplying prior knowledge $\mathsf{K}$ to a NeSy-CBM is equivalent to fixing the inference layer to a corresponding function $\bm{\mathrm{\omega}}^{*}$ (and $\bm{\mathrm{\beta}}^{*}$). Note that whereas $\bm{\mathrm{\omega}}^{*}$ changes based on how reasoning is implemented – e.g., fuzzy vs. probabilistic logic – $\bm{\mathrm{\beta}}^{*}$ does not, as both kinds of reasoning layers behave identically when input any point-mass concept distribution $\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\}$.
Standard assumptions. The maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$, induced respectively by the concept extractor and inference layer, are especially useful for analysis provided the following two standard assumptions about how data are distributed [19, 21, 53]:
Assumption 3.1 (Extrapolability).
@@
This means that the ground-truth concepts $\bm{\mathrm{g}}$ can always be recovered for all inputs $\bm{\mathrm{x}}$ by a sufficiently expressive concept extractor; the $\bm{\mathrm{\alpha}}$ it induces is the identity $\mathrm{id}(\bm{\mathrm{g}}):=\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{g}}\}\in\mathsf{Vert}(\mathcal{A})$.
@@
This ensures that the labels $\bm{\mathrm{y}}$ can be predicted without any ambiguity from the ground-truth concepts $\bm{\mathrm{g}}$. Notice that, not all choices of $\mathcal{G}$ guarantee that these assumptions are met. Small concept spaces may not give a deterministic knowledge, i.e., labels can be confused for one another with only few concepts, or too-arbitrary choices of the constituents may not guarantee their extrapolation from the input, i.e., ground-truth concepts of the input are ambiguous. Nonetheless, both assumptions hold in several NeSy tasks [22] and underlie many works in RSs [19, 63, 20], but formulating a theory that relaxes them is not straightfoward and is technically challenging. In fact, Marconato et al. [19] showed that upon relaxing Assumption˜3.2 it might not be possible to deal with RSs.
The maximum log-likelihood objective is then written as:
@@
Here, $\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}$ is the conditional probability of the ground-truth labels $\bm{\mathrm{y}}$. Notice that under the above assumptions CBMs attaining maximum likelihood perfectly model the ground-truth data distribution $p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})$, see Lemma˜C.1.